(15-2). Unlevered Beta
Counts Accounting’s beta is 1.15 and its tax rate is 40%. If it is financed with 20% debt, what is its unlevered beta?
bL 1.15
Tax rate 40%
wd 20%
Equity ratio 80.00%
bU
Answer #1

Calculate the Unlevered beta as follows:

Unlvered beta(bU) = Beta levered(bL)/(1+((1-tax rate)*(Debt/Equity)))

= 1.15/(1+((1-40%)*(20%/80%)

= 1.15/(1+15%)

= 1.15/1.15

= 1

Therefore unlvered beta is 1.